# Loop Shaping of Hybrid Motion Control with Contact Transition

### v1

@@ Abstract
A standard (stiff) motion control with output displacement feedback cannot handle unforeseen contact with environment without penetrating into soft, i.e. viscoelastic, materials or even damaging brittle or fragile materials. Robotics and mechatronics with tactile and haptic capabilities, and medical assistance systems in particular, place special demands on the advanced motion control systems that should enable safe and harmless contact transitions. This paper demonstrates how the fundamental principles of loop shaping can easily be used to handle the sufficiently stiff motion control with a sensor-free dynamic extension to reconfigure at contact with environment. Hybrid control scheme is proposed. Remarkable feature of the developed approach is that no measurement of the contact force is required and the input signal and measured output displacement are the only quantities used for control design and operation. Experimental scenarios for 1-DOF actuator are shown where the moving tool comes into contact with grape fruits that are soft and penetrable at the same time.
@@ I Introduction
Motion control systems that can come in a well specified or (more importantly) unforseen contact with environmental objects have always been the focus of active research, especially in the field of control and robotics, starting already in the eighties of the last century. For instance, dynamics stability issues during contact with stiff environments were recognized and addressed (often in context of industrial robotics), see [1], when a manipulator in the force control mode comes in touch with a stiff and kinematically constrained environmental object. A milestone was the introduction of the concept of impedance control [2] and later of hybrid impedance control [3], which enabled a deep understanding of the most important (i.e., physically justifiable) interactions and limitations for the controlled impedance–admittance pair of a mechanical motion system in contact with its environment. Motion control of an unconstrained manipulation, on the one hand, and force control of a constrained interaction between the manipulator and its environment, on the other hand, became quickly to ’stumbling block’, especially in view of controlling contact transitions, see e.g. discission with experiments in [4]. For a former brief survey of the force control of manipulators we refer to e.g. [5], while a more recent overview of the force control can be found in the literature like e.g. Springer Handbook of Robotics [6].
@@
The ideas of impedance and admittance control of robotic manipulators, [2, 3], found quickly a way and appreciation in motion control for drives and mechatronic systems [7]. A unified passivity-based control framework for position, torque and impedance control, which use the full state feedback and a dedicated energy shaping with variable gain strategy was also proposed in [8]. Another focus on decomposing (correspondingly switching) the control structure led to hybrid position/force controllers, the first one proposed (most probably) in [9]. A widely adopted strategy of hybrid force/motion control, which aims at controlling the motion along the unconstrained task directions and force along the constrained task directions, is using a certain decomposition which allows simultaneous control of both the contact force and the end-effector motion in two mutually independent subspaces, see [6] for details. Here, a selection strategy for the stiffness/compliance parameters and the desired and feedback variables must be part of the overall control law and is known to be non-trivial and fundamental to the specification of the control task. Although impedance modulation and reconfiguration (corresponding switching) are widely used in hybrid position/force control in robotics, see e.g. [10], equally as in other motion control systems such as hydraulic actuators [11], the problems of transition and stability of structural switching [12] remain among the most relevant. It should be emphasized that during a contact transition, both a hybrid position/force control and the process plant itself undergo a structural change. For sufficiently damped contact transitions and relatively slow dynamics of some available internal state variable experiencing a threshold value upon the contact, a hysteresis relay-based switching policy can be developed, see e.g. [13]. For combining the robustness property of impedance control in stiff contact with the accuracy of admittance control in soft contact various approaches were proposed in robotics. For instance, a predictive instantaneous model impedance control scheme was described in [14], and continuous switching (with duty cycle as design parameter) between controllers with impedance and admittance causality was provided, also with experiments, in [15].
@@
An important causality constraint is that no one motion system can simultaneously impress a force on its environment and impose a displacement or velocity on it. Only one of both control variables, either interaction force or relative motion of the environmental object, can be determined along each degree of freedom [2]. This results from an instantaneous power flow between two or more physical systems. Recall that the power flow is definable as product of an effort and a flow variable, force and velocity for mechanical systems, respectively. Given these basic principles, it seems obvious that different control concepts and an approach for combining them are required to control unconstrained and constrained motion and the resulting contact force.
@@
A frequently appearing question is nevertheless how to handle those manipulator-environment configurations where a control design is a priori constrained by some given structure, the application specification and, most importantly, by the sensors and their arrangement.
@@
In this work, an intuitively understandable (for standard motion control design in frequency domain) approach of reshaping the otherwise stiffly designed feedback controllers is proposed. This way, a smooth and stable contact transition can be guaranteed without applying a more complex control structure. Most importantly, only the measured output displacement is used for feedback and the designed hybrid motion control does not require additional force sensors or observers of internal states. The rest of the paper is organized as follows. In section II we discuss the types of impedance operators that represent an interaction with the environment and relate them to the control loop properties required for a contact transition. Section III introduces the thereupon based hybrid motion control design. A detailed experimental case study with soft but penetrable grape objects coming into contact with the mechanical tool which is driven by the displacement feedback control is demonstrated in section IV. Brief conclusions are drawn in section V.
@@ II Impedance and shaping of disturbance sensitivity function
For motion control at large, with one relative degree of freedom $x$ in the generalized coordinates, one can define the control stiffness, cf. [7], [16], as
@@
Here $f(\cdot)$ is the overall net force (in the generalized coordinates) imposed on the moving body of the motion control system under consideration. For a matched disturbance force $F(t)$ and a feedback control value $u(t)$, which inherently has dimension of a force in motion control, the total net force can be considered as a superposition $f=u+F$. Indeed, when a generic motion control system
$\dot{y}=g(y,u,t),$ (2)
with the sufficiently smooth flow map $g(\cdot)$ and the vector of states variables $y\in\mathbb{R}^{n}$ (often with $n=2$ and $y=(x(t),\dot{x}(t))^{\top}$ as a vector of relative displacement and velocity) is in steady-state, the control should balance the counteracting disturbance. Let us fix next some controlled operation point, in terms of $(u_{0},x_{0})$, and consider this way reduced (1) in Laplace domain. Then, one can write
@@
provided that the Laplace transform of the control stiffness operator exists. Also we note that an equivalence between $\kappa$ and $\tilde{\kappa}$ is valid only for low frequency range, cf. [3]. Recalling that for a linear environment, the impedance is defined as the ratio of the Laplace transforms of effort and flow [17], the mechanical impedance is written as:
$Z(s)=\frac{\hbox{\emph{force}}(s)}{\hbox{\emph{velocity}}(s)},$ (4)
@@
Since a dynamic interaction between two physical bodies implies that one must complement the other, i.e. along any degree for freedom if one is an impedance, the other must be admittance and vice versa, see [2], the condition (5) becomes crucial. It reveals how the motion control system can be designed when an environment is classified by $Z(s)$.
@@
The constitutive equation of the Newtonian fluid in a dashpot, i.e. $F=\alpha\,\dot{x}$ where $\alpha>0$ is viscosity, results in $Z_{v}(s)=\alpha$, meaning the contact environment is resistive, cf. [17]. If one expects the environment to be viscoelastic, i.e. to exhibit also a certain capacitive behavior, then one can assume a Kelvin-Voigt contact, that is schematically shown in Fig. LABEL:fig:environm (b). This leads to the corresponding impedance operator $Z_{ve}(s)=\alpha+\beta/s$. Note that for modeling the environmental impedance, more complex structures than those shown in Fig. LABEL:fig:environm can equally be assumed, and even variable and nonlinear structures might be considered. This would, however, go far beyond the scope of this work, and it turns out to be superfluous for a straightforward design of linear and hybrid motion controllers.
@@
of the motion control system with the input-output plant $G(s)=x(s)(u(s)+F(s))^{-1}$ and feedback controller $C(s)$ which receives the control error $e(s)=r(s)-x(s)$ as input. Obviously, the control reference command $r(s)$ conforms to some application-related specifications, before and after a possible contact with environment and without it. While the reference-to-output transfer function $H(s)=L(s)(1+L(s))^{-1}$ can be determined by shaping $L(s)$, correspondingly designing $C(s)$ to be possibly stiff, i.e. having possibly high bandwidth of $H(s)$ and possibly unity $|H(j\omega)|$ for all angular frequencies $\omega\in(0,+\infty)$, we focus in the following on the disturbance response to $F(s)$. The disturbance-to-output transfer characteristics are given by
@@
and often denoted and disturbance sensitivity function, cf. e.g. [18]. While an ideal position (or velocity) controller should not allow any steady-state or transient deviations for any force imposition on mechanical system, i.e. the controller stiffness should be infinite, cf. [7], a hybrid motion controller with contact transition should allow $S(s)$ to match the $Z(s)$ properties of the contact with environment. Comparing (3), (5), and (7) one can recognize that
@@ III Hybrid motion control
First, we proceed with designing a (sufficiently) stiff feedback motion control, denoted by $C_{s}$. Following the most simple loop shaping methodology and assuming the critically damped dominant pole pair of the closed-loop system, the reference-to-output transfer function yields
@@
Here the control specification is given by only the natural frequency $\omega_{0}$ (approximately to bandwidth) of the closed-loop. Recall that higher $\omega_{0}$ values imply higher stiffness of the controlled system. Using the given system transfer function $G(s)$ and applying the block diagram algebra, the resulting motion control $C_{s}(s)=L(s)G^{-1}(s)$ is
@@
Note that for system plants $G(s)$ with a relative degree $\leq 2$, the control (10) yields a proper transfer function and, thus, can be directly implemented.
@@
that corresponds to a viscous (dashpot-type) contact with environment, cf. section II. Using (7) and the block diagram algebra, the resulting impedance controller, denoted as viscous $C_{v}(s)$, is determined by
$C_{v}(s)=\frac{\alpha s\,G(s)-1}{G}.$ (12)
Note that (12) reveals an improper transfer function so that an additional low-pass filter with a sufficiently high cut-off frequency needs to applied in series with $C_{v}(s)$.
The magnitude response of $G(j\omega)$, $H(j\omega)$, $S_{s}(j\omega)$, and $S_{v}(j\omega)$ transfer functions are depicted in Fig. 2 for the exemplary assumed system plant with two negative real poles at $\{-1000,-10\}$, and the design parameters $\omega_{0}=100$ and $\alpha=100$. Note that both disturbance sensitivity functions are determined according to (7) for each of the feedback controllers $C_{s}$ and $C_{v}$.
Figure 2: Magnitude response of $G(j\omega)$, $G(j\omega)$, $S_{s}(j\omega)$, and $S_{v}(j\omega)$.
It is easy to interpret from the Bode plot of $S_{s}(j\omega)$ that a step-wise disturbance $F(s)$, which appears at the contact instant, will be largely suppressed by the stiff motion control. Following to that, the controlled mechanical motion system will penetrate into the environmental object or moves it away from itself if the latter is unconstrained. Quite the opposite, the viscous impedance control will react to the step-wise contact force $F$ by inducing a repulsive relative displacement in the opposite direction. Here it is worth noting that since it leads to release of the contact, the disturbance becomes zero and the relative motion will stop, cf. below with the experimental results.
@@
If the controlled output displacement is the only measurable state of the motion system (the case we consider in this work), the control reshaping can be triggered exclusively by an information content of the control signal. Assuming some nominal bound of the control signal $U$, that is mostly possible for the given nominal plant $G(s)$, control $C_{s}(s)$, and reference $r(s)$, an overshot $|u(t)|>U$ will indicate appearance of the disturbance force $F$. Note that this strategy can be used in particular when the stiff motion controller $C_{s}$ contains an integral control action. Indeed, during a compensated steady-state motion, an exceeded control force is proportional to an additional disturbing force. Since a dynamic transition from $C_{s}$ to $C_{v}$ should not provoke any undesired transients in direction of the contact with environment, it is worth examining the disturbance-to-control-value transfer characteristics which are given by $U(s)=u(s)/F(s)=C(s)G(s)(1+C(s)G(s))^{-1}$. For both controllers (10) and (12), as exemplary designed above, this is shown by the magnitude response in Fig. 3.
@@
One can recognize that since $r$ is set to zero for $C_{v}$, and the control magnitude response of both $C_{s}$ and $C_{v}$ have nearly the same value, an appearance of the step-wise disturbance $F(t_{c})$ at $t=t_{c}$ will lead to a decrease of $|u(t_{c})|=U$ by the magnitude equal to $|F|$ for $t>t_{c}$. The force imposed on the environmental object under contact drops respectively.
@@ IV-A Motion system and contact scenarios
In the following, we demonstrate an experimental case study of the proposed hybrid motion control with the smooth transition when the moving tool experiences unforeseen contact with environment. The latter is soft yet penetrable, emulating one of the most critical applications of motion control – robotic assistance for medical diagnosis and surgery. For that purpose, a half of the grape is placed in way of the moving mechanical tool, see Fig. LABEL:fig:expsetup, which is controlled by using only the displacement feedback. The relative displacement of the tip of the tool is sensed remotely, i.e. contactless, cf. Fig. LABEL:fig:expsetup. Note that the way how the vertical displacement $x$ is measured provides a relatively high level of the noise and, thus, represents rather a worst case scenario for the control application. The system input and output $v$ (in volt) and $x$ (in meter), respectively, are the only quantities available for the control design and operation.
The experimental motion system in actuated by a voice-coil-motor and has one translational degree of freedom. The rigid mechanical tool is moving in vertical direction and has a relatively low displacement range about 0.015 m. The implemented feedback control is running on the dedicated real-time board with the set sampling rate of 10 kHz. For more technical details, including the physical system parameters, an interested reader is referred to [22].
@@
where $s$ is Laplace variable, and $K$ and $\tau$ are the known system gain and time-constant parameters, respectively. The constant term $D$ constitutes the nominal disturbance due to the gravity force which is known. Therefore, the latter is pre-compensated, so that the system input signal results in
@@
where the feedback controller output $u$ is designed as described in section III.
Due to a free integrator, cf. (13), the identification of the free system parameters was performed in a closed-loop configuration, see [23] for details. The least-squares determined parameter values are $K=0.0408$ and $\tau=0.00668$ sec, while the experimentally measured and identified magnitude response are shown towards each other in Bode diagram in Fig. 5.
@@
Note that an additional electrical time constant of the voice-coil-motor dynamics, which is about $0.0012$ sec, is not explicitly taken into account, equally as not a minor time delay in the input-output loop of the system. Both are neglected in the nominal model (13), cf. [23]. At the same time, they constitute an additional robustness criterion for the motion control system under evaluation and, this way, contribute to a worst-case scenario of the study.

### v2

@@ Abstract
A standard motion control with feedback of the output displacement cannot handle unforeseen contact with environment without penetrating into the soft, i.e. viscoelastic, materials or even damaging the fragile materials. Robotics and mechatronics with tactile and haptic capabilities, and in particular medical robotics for example, place special demands on the advanced motion control systems that should enable the safe and harmless contact transitions. This paper shows how the basic principles of loop shaping can be easily used to handle sufficiently stiff motion control in such a way that it is extended by sensor-free dynamic reconfiguration upon contact with the environment. A thereupon based hybrid control scheme is proposed. A remarkable feature of the developed approach is that no measurement of the contact force is required and the input signal and the measured output displacement are the only quantities used for design and operation. Experiments on 1-DOF actuator are shown, where the moving tool comes into contact with grapes that are soft and simultaneously penetrable.
@@ I Introduction
Motion control systems that can come into precisely defined or (more importantly) unpredictable contact with objects in the environment have always been the focus of active research, especially in the field of control and robotics, and have done so since the eighties of the last century. For instance, the dynamics stability issues during a contact with stiff environments were recognized and addressed (often in context of industrial robotics), see [1], when a manipulator in the force control mode comes in touch with a stiff and kinematically constrained environmental object. A milestone was the introduction of the concept of impedance control [2] and later of hybrid impedance control [3], which enabled a deep understanding of the most important (i.e., physically justifiable) interactions and constraints for the controlled impedance-admittance pair of a mechanical motion system in contact with its environment. Motion control of an unconstrained manipulation, on the one hand, and force control of a constrained interaction between the manipulator and its environment, on the other hand, became quickly to ’stumbling block’, especially in view of controlling contact transitions, see e.g. discussion with experiments in [4]. For a former compact survey of the force control of manipulators we refer to e.g. to the work [5], while a more recent overview of the force control can be found in the robotic literature like e.g. Springer Handbook of Robotics [6].
@@
The ideas of impedance and admittance control of robotic manipulators, [2, 3], found quickly a way and appreciation in motion control for drives and mechatronic systems [7]. A unified passivity-based control framework for position, torque and impedance control, which uses the full state feedback and a dedicated energy shaping with variable gain strategy was also proposed in [8]. Another focus on decomposing (correspondingly switching) the control structure led to hybrid position/force controllers, the first one (most probably) proposed in [9]. A widely adopted strategy of hybrid force/motion control, which aims at controlling the motion along the unconstrained task directions and the force along the constrained task directions, is using a certain decomposition which allows simultaneous control of both the contact force and end-effector motion in two mutually independent subspaces, see [6] for details. Here, a selection strategy for the stiffness/compliance parameters and the desired and feeded back variables must be part of the overall control law and is known to be non-trivial and fundamental for the specification of the control task. Although impedance modulation and reconfiguration (correspondingly switching) are widely used in the hybrid position/force control in robotics, see e.g. [10], equally as in other motion control systems such as hydraulic actuators [11], the problems of transition and stability of the structural switching [12] remain among the most relevant. It should be emphasized that during a contact transition, both a hybrid position/force control and the process plant itself undergo a structural change. For sufficiently damped contact transitions and a relatively slow dynamics of an available internal state variable that experiences a threshold value upon contact, a switching strategy based on hysteresis relays can be applied, see e.g. [13]. For combining the robustness property of an impedance control in the stiff contacts with the accuracy of an admittance control in the soft contacts, various approaches were proposed in robotics. For instance, a predictive instantaneous model impedance control scheme was described in [14], and a continuous switching (with duty cycle as design parameter) between the controllers with impedance and admittance causality was provided (also with experiments) in [15].
@@
An important causality constraint is that no one motion system can simultaneously impress a force on its environment and impose a displacement or velocity on it. Only one of both control variables, either interaction force or relative motion of the environmental object, can be determined along each degree of freedom [2]. This results from an instantaneous power flow between two or more physical systems. Recall that the power flow is definable as product of an effort and a flow variable, the force and velocity for mechanical systems, respectively. Given these basic principles, it seems obvious that different control concepts and an approach for combining them are required to control the unconstrained and constrained motion and the resulting contact force.
@@
A frequently appearing question is nevertheless how to handle those manipulator-environment configurations where the control design is a-priori constrained by some given structure and/or application specification and, most importantly, by the available sensors and their arrangement.
@@
In this work, an intuitively understandable (for standard control design in frequency domain) approach of reshaping the otherwise stiffly designed feedback controllers is proposed. This way, a smooth and stable contact transition can be guaranteed without applying a more complex control structure. Most importantly, only the measured output displacement is used for feedback and the designed hybrid motion control does not require additional force sensors or observers of internal states. The rest of the paper is organized as follows. In section II, we discuss the types of impedance operators that represent an interaction with environment and relate them to the control loop properties required for a contact transition. Section III introduces the thereupon based hybrid motion control design. A detailed experimental case study with soft but penetrable grape objects coming into contact with the feedback controlled mechanical tool is provided in section IV. Brief conclusions are in section V.
@@ II Impedance and shaping of disturbance sensitivity function
For motion control at large, with one relative degree of freedom $x$ specified in the generalized coordinates, one can define the control stiffness, cf. [7], [16], as
@@
Here $f(\cdot)$ is the overall net force (in the generalized coordinates) imposed on the moving body of the motion control system under consideration. For a matched disturbance force $F(t)$ and a feedback control value $u(t)$, which inherently has dimension of force in the motion control, the total net force can be considered as a superposition $f=u+F$. Indeed, when a generic motion control system
$\dot{y}=g(y,u,t)$ (2)
with the sufficiently smooth flow map $g(\cdot)$ and the vector of states variables $y\in\mathbb{R}^{n}$ (often a vector of relative displacements and velocities) is in the steady-state, the control should balance the counteracting disturbance. Let us next fix some controlled operation point, in terms of $(u_{0},x_{0})$, and consider this way reduced (1) in the Laplace domain. Then, one can write
@@
provided that the Laplace transform of the control stiffness operator exists. Also we note that an equivalence between $\kappa$ and $\tilde{\kappa}$ is valid only for low frequency range, cf. [3]. Recalling that for a linear environment, an impedance is defined as the ratio of the Laplace transforms of effort and flow [17], the mechanical impedance is (symbolically) written as
$Z(s)=\frac{\hbox{force}(s)}{\hbox{velocity}(s)},$ (4)
@@
Since a dynamic interaction between two physical bodies implies that one must complement the other – i.e. along any degree for freedom if one is an impedance then the other must be admittance and vice versa [2] – the condition (5) becomes crucial. It reveals how the motion control system can be designed when an environment is classified by $Z(s)$.
@@
The constitutive equation of the Newtonian fluid in a dashpot, i.e. $F=\alpha\,\dot{x}$ where $\alpha>0$ is viscosity, results in $Z_{v}(s)=\alpha$, meaning the contact environment is resistive, cf. [17]. If one expects the environment to be viscoelastic, i.e. to exhibit also a certain capacitive behavior, then one can assume a Kelvin-Voigt contact, that is schematically shown in Fig. LABEL:fig:environm (b). This leads to the corresponding impedance operator $Z_{ve}(s)=\alpha+\beta/s$. Note that for modeling the environmental impedance, more complex structures than those shown in Fig. LABEL:fig:environm can equally be assumed, and even variable and nonlinear structures might be considered. This would, however, go far beyond the scope of this work, and it turns out to be superfluous for a straightforward design of the linear and hybrid motion controllers.
@@
of the motion control system with the input-output plant $G(s)=x(s)(u(s)+F(s))^{-1}$ and feedback controller $C(s)$. The latter receives the control error $e(s)=r(s)-x(s)$ as input. Obviously, the control reference command $r(s)$ conforms to some application-related specifications before and after a possible contact with environment and without it. While the reference-to-output transfer function $H(s)=L(s)(1+L(s))^{-1}$ can be determined by shaping $L(s)$, correspondingly by designing $C(s)$ to be possibly stiff, i.e. having possibly high bandwidth of $H(s)$ and possibly unity $|H(j\omega)|$ for all angular frequencies $\omega\in(0,+\infty)$, we focus on the disturbance response to $F(s)$ in the following. The disturbance-to-output transfer characteristics are given by
@@
and often denoted as disturbance sensitivity function, cf. e.g. [18]. While an ideal position (or velocity) controller should not allow any steady-state or transient deviations for any force imposition on the mechanical system, i.e. the controller stiffness should be infinite, cf. [7], a hybrid motion controller with contact transition should allow $S(s)$ to match the $Z(s)$ properties of the contact with environment. Comparing (3), (5), and (7) one can recognize that
@@ III Hybrid motion control
First, we proceed with designing a (sufficiently) stiff feedback motion control denoted by $C_{s}$. Following the most simple loop shaping methodology and assuming the critically damped dominant pole pair of the closed-loop control system, the reference-to-output transfer function yields
@@
The control specification is then given by only the natural frequency $\omega_{0}$ (approximately equal to the bandwidth) of the closed-loop. Recall that higher $\omega_{0}$ values imply higher stiffness of the controlled system. Using the given system transfer function $G(s)$ and applying the block diagram algebra, the resulting motion control $C_{s}(s)=L(s)G^{-1}(s)$ is
@@
Note that for the system plants $G(s)$ with a relative degree $\leq 2$, the control (10) yields a proper transfer function and, thus, can be directly implemented.
@@
that corresponds to a viscous (dashpot-type) contact with environment, cf. section II. Using (7) and the block diagram algebra, the resulting impedance controller (denoted further as viscous) is determined by
$C_{v}(s)=\frac{\alpha s\,G(s)-1}{G(s)}.$ (12)
Note that (12) reveals an improper transfer function so that an additional low-pass filter with a sufficiently high cut-off frequency needs to be applied in series with $C_{v}(s)$.
The magnitude response of $G(j\omega)$, $H(j\omega)$, $S_{s}(j\omega)$, and $S_{v}(j\omega)$ transfer functions are depicted in Fig. 2 for an exemplary assumed system plant with two negative real poles at $\{-1000,-10\}$, and the design parameters $\omega_{0}=100$ and $\alpha=100$. Note that both disturbance sensitivity functions are determined according to (7) for each of the feedback controllers $C_{s}$ and $C_{v}$.
Figure 2: Magnitude response of $G(j\omega)$, $H(j\omega)$, $S_{s}(j\omega)$, and $S_{v}(j\omega)$.
It is easy to interpret from the Bode plot of $S_{s}(j\omega)$ that a step-wise disturbance $F(s)$, which appears at the contact instant, will be largely suppressed by the stiff motion control. Following to that, the controlled mechanical motion system will continue to move and penetrate into the environmental object, or moves it away from itself if the latter is not fixed. Quite the opposite, the viscous impedance control will react to the step-wise contact force $F$ by inducing a repulsive relative displacement in the opposite direction. Here it is worth noting that since it leads to release from the contact, the disturbance becomes zero and the relative motion will stop, cf. below with the experiments.
@@
If the controlled output displacement is the only measurable state of a motion system (the case that we also consider in this work), the control reshaping can be triggered exclusively by an information content of the control signal. Assuming some nominal bound of the control signal $U$, that is mostly possible for the given nominal plant $G(s)$, control $C_{s}(s)$, and reference $r(s)$, an overshot $|u(t)|>U$ will indicate the appearance of a disturbance force $F$. Note that this strategy can be used in particular when the stiff motion controller $C_{s}$ contains an integral control action. Indeed, during a compensated steady-state motion, an exceeded control force is proportional to an additional disturbing force. Since a dynamic transition from $C_{s}$ to $C_{v}$ should not provoke any undesired transients in direction of the contact with environment, it is worth examining the disturbance-to-control-value transfer characteristics which are given by $U(s)=u(s)/F(s)=C(s)G(s)(1+C(s)G(s))^{-1}$. For both controllers (10) and (12), as designed exemplary above, this is shown by the magnitude response in Fig. 3.
@@
One can recognize that since $r$ is set to zero for $C_{v}$, and the control magnitude response of both $C_{s}$ and $C_{v}$ have nearly the same value, an appearance of the step-wise disturbance $F(t_{c})$ at $t=t_{c}$ will lead to a decrease of $|u(t_{c})|=U$ by the magnitude equal to $|F|$ for $t>t_{c}$. Then, the force imposed on the environmental object under contact drops respectively.
@@ IV-A Motion system and contact scenarios
In the following, we demonstrate an experimental case study of the proposed hybrid motion control with a smooth transition when the moving tool experiences unforeseen contact with environment. The latter is soft yet penetrable, emulating one of the most critical applications of the motion control – robotic assistance for medical diagnosis and surgery. For that purpose, a half of the grape is placed in the way of the moving mechanical tool, see Fig. LABEL:fig:expsetup, which is controlled by using only the displacement feedback. The relative displacement of the tip of the tool is sensed remotely, i.e. contactless, cf. Fig. LABEL:fig:expsetup. Note that the way how the vertical displacement $x$ is measured provides a relatively high level of the noise and, thus, represents rather a worst case scenario for the control application. The system input $v$ (in volt) and output $x$ (in meter) are the only quantities available for the control design and operation. The experimental motion system in actuated by a voice-coil-motor and has one translational degree of freedom. The rigid mechanical tool is moving in the vertical direction and has a relatively low displacement range about 0.015 m. The implemented feedback control is running on the dedicated real-time board with the set sampling rate of 10 kHz. For more technical details, including the physical system parameters, an interested reader is referred to [22].
@@
where $s$ is the Laplace variable, and $K$ and $\tau$ are the known system gain and time-constant parameters, respectively. The constant term $D$ constitutes the nominal disturbance due to the gravity force which is known. Therefore, the latter is pre-compensated, so that the system input signal results in
@@
where the feedback controller output $u$ is designed as described above in section III.
Due to a free integrator, cf. (13), the identification of the free system parameters was performed in a closed-loop configuration, see [23] for details. The least-squares determined parameter values are $K=0.0408$ and $\tau=0.00668$ sec, while the experimentally measured and identified magnitude response are shown over each other in the Bode diagram in Fig. 5.
@@
Note that an additional electrical time constant of the voice-coil-motor dynamics, which is about $0.0012$ sec, is not explicitly taken into account, equally as not a minor time delay in the input-output loop of the system. Both are neglected in the nominal model (13), cf. [23]. At the same time, they constitute an additional robustness criterion for the motion control system under evaluation and, this way, contribute to a worst-case scenario under evaluation.